The cross-sectional area of the aorta A0 (the major blood vessel emerging from the heart) of normal resting person is 3 cm2, and the speed v0 of the blood through it is 30 cm/s. A typical capillary (with diameter approximately 6×10-6 m) has a flow speed v of 0,05 cm/s. How many capillaries does such a person have?
1
Expert's answer
2020-12-25T14:06:18-0500
"A_0v_0=0.25\pi d^2nv\\(3)(30)=0.25\pi (6\cdot10^{-6})^2(0.05)n\\n=6.37\cdot10^{13}"